OpenStudy (anonymous):

How long it will take for $300 to double when invested at 6% annual interest compounded twice a year

OpenStudy (kropot72):

600 = 300(1 + 0.03)^x = 300 * 1.03^x 1.03^x = 600/300 = 2 x * log 1.03 = log 2 x = (log 2)/(log 1.03) = 23.45 six month periods It will take approximately 11 1/2 years to double the $300